\begin{thebibliography}{99}

%\bibitem{Amitsur_1980} S.A. Amitsur, {\it On the characteristic polynomial of a sum of matrices}, Linear and Multilinear Algebra {\bf 8} (1980), 177--182.

%\bibitem{Aslaksen95} H. Aslaksen, E.-C. Tan, C.-B. Zho, {\it Invariant theory of special orthogonal groups}, Pac. J. Math. {\bf 168} (1995), No.~2, 207--215.

%\bibitem{ADS_2006}H. Aslaksen, V. Drensky, L. Sadikova, {\it Defining re\-lations of %invariants of two $3\times 3$ matrices}, J. Algebra {\bf 298} (2006), 41--57.

\bibitem{Cavalcante_Lopatin_1} F.B. Cavalcante, A. Lopatin, {\it Separating invariants of three nilpotent $3\times3$ matrices}, Linear Algebra and its Applications {\bf 607} (2020), 9--28. 

%\bibitem{DW00}H. Derksen, J. Weyman, {\it Semi-invariants of qui\-vers 
%and saturation for Little\-wood-Richardson coeffi\-ci\-ents}, J. Amer. Math. Soc. {\bf
%13} (2000), 467--479.

%\bibitem{DW_LR_02}H. Derksen, J. Weyman, {\it On the Littlewood-Ri\-chard\-son
%polynomials}, J. Algebra {\bf 255} (2002), No.~2, 247--257.

\bibitem{DerksenKemper_book}H. Derksen and G. Kemper, {\it Computational Invariant Theory}, Invariant Theory and Algebraic Transformation Groups, I. Encyclopaedia of Mathematical Sciences, 130. Springer-Verlag, Berlin, 2002. x+268 pp.

%\bibitem{DW_GenQuiv_2002}H. Derksen, J. Weyman, {\it Generalized quivers asso\-ciated to reductive groups}, Colloq. Math. {\bf 94} (2002), No.~2., 151--173.

%\bibitem{Domokos98}M. Domokos, {\it Invariants of quivers and wreath pro\-ducts},
%Comm. Algebra {\bf 26} (1998), 2807--2819.

%\bibitem{DZ01}M. Domokos, A.N. Zubkov,
%{\it Semi-invariants of quivers as determinants}, Transform. Groups {\bf 6} (2001),
%No.~1, 9--24.

%\bibitem{Domokos_gen_2002}M. Domokos, {\it Finite generating system of matrix invariants},
%Math. Pannon. {\bf 13} (2002), No.~2, 175--181.

\bibitem{Domokos_2007} M.  Domokos,  {\it Typical  separating  invariants},  Transform.   Groups  {\bf 12} (2007), 49--63.

%\bibitem{Djokovic_O4}D.$\rm\check{Z}$. \DJ okovi$\rm\acute{c}$ and M.L. MacDonald, {\it Orthogonal invariants of a matrix of order four and applications}, J. Pure Appl. Algebra {\bf 202} (2005), 259--283.
%

%\bibitem{Djokovic_Osmall}D.$\rm\check{Z}$. \DJ okovi$\rm\acute{c}$, {\it On orthogonal and special orthogonal invariants of a single matrix of small order}, Linear and Multilinear Algebra {\bf 57} (2009), no.~4, 345--354.


\bibitem{DKZ_2002}M. Domokos, S.G. Kuzmin, A.N. Zubkov, {\it Rings of matrix invariants in positive characteristic}, J. Pure Appl. Algebra {\bf 176} (2002), 61--80.  %

%\bibitem{Domokos_Frenkel} M. Domokos, P.E.~Frenkel, {\it Mod $2$ indecomposable orthogonal
%invariants}, Adv. Math. {\bf 192} (2005), 209--217.

%\bibitem{Domokos_2016}M. Domokos, {\it Degree bound for separating invariants of abelian groups}, Proc. of Amer. Math. Soc., {\bf 145} (2016),  3695-3708. 

 

\bibitem{Domokos20} M. Domokos, {\it Characteristic free description of semi-invariants of $2\times2$ matrices}, J. Pure Appl. Algebra {\bf 224} (2020), no. 5, 106220. 
%https://doi.org/10.1016/j.jpaa.2019.106220.

\bibitem{Domokos20Add} M. Domokos, {\it Addendum to ``Characteristic free description of semi-invariants of $2\times2$ matrices'' [J. Pure Appl. Algebra 224 (2020), no. 5,  106220]}, J. Pure Appl. Algebra {\bf 224} (2020), no. 6, 106270. %https://doi.org/10.1016/j.jpaa.2019.106270.

%\bibitem{Donkin85}S. Donkin, {\it Rational representations of algebraic groups:
%tensor products and filtrations}, Lecture Notes in Math. {\bf 1140} (1985), Springer,
%Berlin, Heidelberg, New York.

%\bibitem{Donkin90}S. Donkin, {\it The normality of closures of conjugacy classes of matrices}, Invent. Math. {\bf 101} (1990), 717--736.

\bibitem{Donkin92a}S. Donkin, {\it Invariants of several matrices}, Invent. Math. {\bf 110} (1992), 389--401.

%\bibitem{Donkin93a}S. Donkin, {\it Invariant functions on matrices}, Invent. Math. {\bf 113} (1993), 23--43.

%\bibitem{Donkin93}S. Donkin, {\it On tilting modules for algebraic groups}, Math. Z. {\bf 212} (1993), 39--60.

%\bibitem{Donkin94}S. Donkin, {\it Polynomial invariants of representations of quivers}, Comment. Math. Helvetici {\bf 69} (1994), 137--141.

%\bibitem{Drensky_survey_2007}V. Drensky, {\it Computing with matrix invariants}, Math. %Balkanica (N.S.) {\bf 21} (2007), No.~1-2, 141--172. 

%\bibitem{Drensky_Sadikova_4x4}V. Drensky, L. Sadikova, {\it Generators of invariants of two $4\times 4$ matrices}, C. R. Acad. Bulgare Sci. {\bf 59} (2006), No.~5, 477--484.

\bibitem{DraismaKemperWehlau_2008}J. Draisma, G. Kemper, D. Wehlau, {\it Polarization of separating invariants}, Canad. J.
Math. {\bf 60}, (2008) No.~3, 556--571.

%\bibitem{Dufresne_2009}E. Dufresne, {\it Separating invariants and finite reflection groups}, Advances in Mathematics {\bf 221} (2009), 1979--1989.

%\bibitem{Dufresne_Elmer_Kohls_2009}E. Dufresne, J. Elmer, M. Kohls, {\it The Cohen-Macaulay property of separating invariants of finite groups}, Transformation Groups {\bf 14} (2009),  771--785.

%\bibitem{Dufresne_Elmer_Sezer_2013}E. Dufresne, J. Elmer, M. Sezer, {\it Separating invariants for arbitrary linear actions of the additive group}, Manuscripta Mathematica {\bf 143}, (2013), No. 1--2, 207--219.

%\bibitem{Dufresne_Kohls_2013}E. Dufresne, M. Kohls,  {\it The separating variety for the basic representations of the additive group}, J. Algebra {\bf 377} (2013), 269--280.

%\bibitem{Dufresne_2015}E. Dufresne, {\it Separating invariants and local cohomology}, Advances in Mathematics {\bf 270} (2015), 565--581.

%\bibitem{Elmer_Kohls_2012}J. Elmer, M. Kohls, {\it  Separating invariants for the basic $\mathbb{G}_a$-actions},  Proc. Amer. Math. Soc. {\bf 140} (2012), 135-146.

%\bibitem{Elmer_Kohls_2014}J. Elmer, M. Kohls,  {\it Zero-separating invariants for finite groups}, Journal of Algebra  {\bf 411} (2014), 92--113.

%\bibitem{Elmer_Kohls_2016}J. Elmer, M. Kohls,  {\it Zero-separating invariants for linear algebraic groups}, Proc. Edinburgh  Math. Soc. {\bf 59} (2016), 911--924. 

%\bibitem{Lopatin_O_separ}R.J.S. Ferreira, A.A. Lopatin,  {\it Separating orthogonal invariants in low dimension case}, submitted.

%\bibitem{Jing_2015}N. Jing, {\it Unitary and orthogonal equivalence of sets of matrices}, Linear Algebra and its Applications {\bf 481} (2015), 235--242.

%\bibitem{Kirichenko}Yu.P. Drozd, V.V. Kirichenko,  {\it Finite dimensional algebras}, Springer-Verlag, 1994.

%\bibitem{Formanek_1987} E. Formanek, {\it The invariants of $n\times n$ matrices},mLecture Notes in Math. {\bf 1278} (1987), 18--43.

%\bibitem{Formanek_1991} E. Formanek, {\it The polynomial identities and,in\-va\-riants of $n\times n$ matrices}, Regional Conference series in Mathematics {\bf 78}, Providence, RI; American Math. Soc., 1991.

%\bibitem{Gabriel72}P. Gabriel, {\it Unzerlegbare Darstellungen I}, Manuscr. Math. {\bf 6} (1972), 71--103.

%\bibitem{Gantmacher}F.R. Gantmacher, {\it The theory of matrices, Volume~2}, AMS Chelsea Publishing, 1959, 276 pp.

%\bibitem{Grosshans97}F.D. Grosshans, {\it Algebraic homogeneous spaces and invariant theory}, Lecture Notes in Math. {\bf 1673} (1997), Springer: Berlin, Heidelberg, New York.

%\bibitem{Hashimoto_2001} M. Hashimoto, {\it Good filtrations of symmetric algebras and strong F-regularity of invariant subrings}, Math. Z. {\bf 236} (2001), 605--623.

%\bibitem{Hilbert_old}D. Hilbert, {\it \"{U}ber die vollen Invariantensysteme}, Math. Ann. {\bf 42} (1893), 313--373.

%\bibitem{Hilbert_new}D. Hilbert, {\it \"{U}ber die vollen Invariantensysteme}, Ges. Abh., II, Springer--Verlag, 1970, 287--344.

%\bibitem{Hodge_Pedoe_I} W.V.D. Hodge, D. Pedoe, {\it Methods of algebraic geometry}, Vol. I, Cambridge, Univ. Press, 1947.

%\bibitem{Hodge_Pedoe_II} W.V.D. Hodge, D. Pedoe, {\it Methods of algebraic geometry}, Vol. II, Cambridge, Univ. Press, 1952.

%\bibitem{Kraft}H. Kraft, {\it Geometrische Methoden in der Invariantentheorie}, Friedr. Vieweg \& Sohn: Braunschweig, Wiesbaden, 1985.

%\bibitem{Kohls_Kraft_2010} M.  Kohls, H. Kraft,  {\it Degree bounds for separating invariants}, Math. Research Letters, {\bf 17} (2010) 10001--10012.

%\bibitem{Kohls_Sezer_2013}M. Kohls,  M. Sezer, {\it Separating invariants for the Klein group and cyclic groups}, International Journal of Mathematics, {\bf 24} (2013),  No. 6, 1350046. 

\bibitem{Lopatin_separating2x2} I. Kaygorodov, A. Lopatin, Yu. Popov, {\it Separating invariants for $2\times 2$ matrices}, Linear Algebra and its Applications {\bf 559} (2018), 114--124. 

\bibitem{Kemper_Lopatin_Reimers_2} G. Kemper, A. Lopatin, F. Reimers, {\it Separating invariants over finite fields}, Journal of Pure and Applied Algebra {\bf 226} (2022), 106904. 

 %\bibitem{Le_Bruyn_Procesi_90}L. Le Bruyn, C. Procesi, {\it Semi-simple repre\-sen\-ta\-tions of quivers}, Trans. Amer. Math. Soc. {\bf 317} (1990), 585--598.

\bibitem{Lopatin_Sib}A.A. Lopatin, {\it The invariant ring of triples of $3\times 3$ matrices over a field of arbitrary characteristic}, Sibirsk. Mat. Zh. {\bf 45} (2004), No.~3, 624--633 (Russian). English translation: Siberian Mathematical Journal {\bf 45} (2004), No.~3, 513--521.

\bibitem{Lopatin_Comm1}A.A. Lopatin, {\it The algebra of invariants of $3\times 3$ matrices over a field of arbitrary characteristic}, Commun. Algebra {\bf 32} (2004), no.~7, 2863--2883.

\bibitem{Lopatin_Comm2}A.A. Lopatin, {\it Relatively free algebras with the identity $x^3=0$}, Commun. Algebra {\bf 33} (2005), no.~10, 3583--3605.

%\bibitem{Lopatin_bplp}A.A. Lopatin, {\it On block partial linearizations of the pfaffian}, %Linear Algebra Appl. {\bf 426/1} (2007), 109--129.

%\bibitem{LZ1}A.A. Lopatin, A.N. Zubkov,
%{\it Semi-invariants of mixed representations of quivers}, Transform. Groups {\bf 12} (2007), N2, 341--369.

%\bibitem{LZ_survey}A.A. Lopatin, A.N. Zubkov, {\it Representations of quivers, their %generalizations and invariants}, Herald of Omsk Univ., Special Issue: Combinatorial Methods %of Algebra and Complexity of Computations (2008), 9--24.

\bibitem{Lopatin_so_inv}A.A. Lopatin, {\it Invariants of quivers under the ac\-tion of classical groups}, J. Algebra {\bf 321} (2009), 1079--1106.

%\bibitem{Lopatin_2222}A.A. Lopatin, {\it Indecomposable invariants of quivers for dimension %$(2,\ldots,2)$ and maximal paths}, submitted, arXiv: 0704.2411.

%\bibitem{Lopatin_semi2222}A.A. Lopatin, {\it Minimal generating set for semi-invariants of quivers of dimension two}, Linear Algebra and its Applications, {\bf 434} (2011), no.~8, 1920--1944.

\bibitem{Lopatin_Oskew}A.A. Lopatin, {\it Orthogonal invariants of skew-symmetric matrices}, Linear and Multilinear Algebra {\bf 59} (2011), 851--862.

%\bibitem{Lopatin_O3}A.A. Lopatin, {\it On minimal generating systems for matrix $O(3)$-invariants}, Linear and Multilinear Algebra {\bf 59}, (2011) 87--99.

\bibitem{Lopatin_Orel}A.A. Lopatin, {\it Relations between $O(n)$-invariants of several matrices}, Algebras and Representation Theory {\bf 15} (2012), 855--882.

\bibitem{Lopatin_Ofree}A.A. Lopatin, {\it Free relations for matrix invariants in the modular case}, Journal of Pure and Applied Algebra {\bf 216} (2012), 427--437.

%\bibitem{Lopatin_xn}A.A. Lopatin, {\it On the nilpotency degree of the algebra with identity $x^n=0$}, Journal of Algebra {\bf 371} (2012), 350--366.

\bibitem{Lopatin_mgs_skewO4}A.A. Lopatin, {\it Minimal system of generators for $O(4)$-invariants of two skew-symmetric matrices}, Linear and Multilinear Algebra, {\bf 66} (2018), no.~2, 347--356.

%\bibitem{Lopatin_Reimers_1} A. Lopatin, F. Reimers, {\it Separating invariants for multisymmetric polynomials}, Proceedings of Amer. Math. Soc. {\bf 149} (2021),  497--508. 

\bibitem{Lopatin_IndOrt}A.A. Lopatin, {\it Indecomposable orthogonal invariants of several matrices over a field of positive characteristic}, International Journal of Algebra and Computation {\bf 21} (2021), no.~1, 161--171.

%\bibitem{Mathieu90}O. Mathieu, {\it Filtrations of $G$-modules},
%Ann. Scient. Ec. Norm. Sup. (2) {\bf 23} (1990), 625--644.

%\bibitem{Nakamoto_2002}K. Nakamoto, {\it The structure of the invariant ring of two %matrices of degree 3}, J. Pure Appl. Algebra {\bf 166} (2002), 125--148.

%\bibitem{Neusel_2009} M.D. Neusel, M. Sezer, {\it Separating invariants for modular $p$-groups and groups acting diagonally}, Math. Res. Lett. {\bf 16} (2009),  No.6, 1029--1036.

\bibitem{Procesi76}C. Procesi, {\it The invariant theory of $n\times n$ matrices}, Adv. Math. {\bf 19} (1976), 306--381.

%\bibitem{Procesi_1984}C. Procesi, {\it Computing with $2\times 2$ matrices}, J. Algeb\-ra {\bf 87} (1984), 342--359.


\bibitem{Razmyslov74}Yu.P. Razmyslov, {\it Trace identities of full matrix algebras over a field of characteristic $0$}, Izv. Akad. Nauk SSSR Ser. Mat. {\bf 38} (1974), No.~4, 723--756 (Russian). English translation: Math. USSR Izv. {\bf 8} (1974), No.~4, 727--760.

%\bibitem{Sezer_2009}M. Sezer, {\it Constructing modular separating invariants}, Journal of Algebra, {\bf 322} (2009),  No. 11, 4099--4104.


%\bibitem{Sezer_2011}M. Sezer, {\it Explicit separating invariants for cyclic $P$-groups}, Journal of Combinatorial Theory, Series A, {\bf 118} (2011), 681--689.


%\bibitem{Shmelkin}D.A. Shmelkin, {\it Signed quivers, symmetric quivers,
%and root systems}, J. Lond. Math. Soc., II. Ser. {\bf 73} (2006), No.~3, 586--606.

%\bibitem{Sibirskii_1967}K.S. Sibirskii, {\it Unitary and orthogonal invariants of %matrices}, Dokl. Akad. Nauk SSSR {\bf 172} (1967), 40--43 (Russian).
%English translation: Soviet Math. Dokl. {\bf 8} (1967), 36--40.

%\bibitem{Sibirskii_1967_2x2}K.S. Sibirskii, {\it Orthogonal invariants of the system of %$2\times2$ matrices},
%Mat. Issled. {\bf 2} (1967), vyp. 4, 124--135 (Russian).

\bibitem{Sibirskii_1968}K.S. Sibirskii, {\it Algebraic invariants of a system of matrices}, Sibirsk. Mat. Zh. {\bf 9} (1968), No.~1, 152--164 (Russian). English translation: Soviet Math. Dokl. {\bf 8} (1967), 36--40.

%\bibitem{Schofield_van_den_Bergh_01}A. Schofield, M. van den Bergh, {\it Semi-invariants
%of quivers for arbitrary dimension vectors}, Indag. Math. (N.S.) {\bf 12} (2001), No.~1,
%125--138.

\bibitem{Spencer1} A.J.M. Spencer, R.S. Rivlin, {\it The theory of matrix polynomials and its application to the mechanics of isotropic continua}, Arch. Rational Mech. Anal. {\bf 2} (1958/1959), 309--336. 

\bibitem{Spencer2} A.J.M. Spencer, R.S. Rivlin, {\it Finite integrity bases for five or fewer symmetric $3\times3$ matrices}, Arch. Rational Mech. Anal. {\bf 2} (1958/1959), 435--446.

\bibitem{Spencer3} A.J.M. Spencer, {\it Further results in the theory of matrix polynomials}, Arch. Rational Mech. Anal. {\bf 4} (1960), 214--230.

\bibitem{Spencer4} A.J.M. Spencer, {\it The invariants of six symmetric $3\times3$ matrices}, Arch. Rational Mech. Anal. {\bf 7} (1961), 64--77.

\bibitem{Spencer_book} A.J.M. Spencer, {\it Theory of invariants}, Continuum Physics, Vol. I, Part III, Academic Press, New York, 1971.

%\bibitem{Teranishi_1986}Y. Teranishi, {\it The ring of invariants of matrices}, Nagoya Math. J. {\bf 104} (1986), 149--161.

%\bibitem{Teranishi_1988}Y. Teranishi, {\it The Hilbert series of rings of matrix concomitants}, Nagoya Math. J. {\bf 111} (1988), 143--156.

%\bibitem{Wellstein_1930}J. Wellstein, {\it \"Uber symmetrische, alternierende und orthogonale Normalformen von Matrizen}, J. reine angew. Math. {\bf 163} (1930), 166--182.

%\bibitem{Weyl46} H. Weyl, {\it The classical groups --- their invariants and
%representations},  Princeton Univ. Press, Prin\-ceton, N.J., 1946.

\bibitem{Zubkov96}A.N. Zubkov, {\it On a generalization of the Razmyslov--Procesi theorem}, Algebra and Logic {\bf 35} (1996), No.~4, 241--254.

%\bibitem{Zubkov_preprint} A.N. Zubkov, {\it  Mixed representations of quivers  and relative problems}, Bielefeld University, SFB {\bf 343} (2000), Preprint 00-094.

\bibitem{Zubkov99}A.N. Zubkov, {\it Invariants of an adjoint action of classical groups}, Algebra and Logic {\bf 38} (1999), No.~5, 299--318.

%\bibitem{Zubkov_Fund_Math_01}A.N. Zubkov, {\it The Razmyslov-Procesi theorem for quiver representations}, Fundam. Prikl. Mat. {\bf 7} (2001), No.~2, 387--421 (Russian).

\bibitem{ZubkovI}A.N. Zubkov, {\it Invariants of mixed representations of quivers I}, J. Algebra Appl. {\bf 4} (2005), No.~3, 245--285.

%\bibitem{ZubkovII}A.N. Zubkov,  {\it Invariants of mixed
%representations of quivers II: Defining relations and applications}, J. Algebra Appl. {\bf 4} (2005), No.~3, 287--312.

\end{thebibliography}

\documentclass[11pt, twoside, reqno]{amsart}
\usepackage{xcolor, amstext, amsfonts, amssymb, amsbsy, latexsym}
\usepackage{enumerate}
\usepackage[T1]{fontenc}
\usepackage{xy, hhline}
\newcounter{num}[section] %

\newenvironment{theo}
{\refstepcounter{num}%
\bigskip\noindent{\bf Theorem~\arabic{section}.\arabic{num}. }\it}
{\smallskip}

\newenvironment{lemma}
{\refstepcounter{num}%
\bigskip\noindent{\bf Lemma~\arabic{section}.\arabic{num}. }\it}
\newenvironment{eq}{\begin{equation}}{\end{equation}}

\newcommand{\si}{\sigma}
\newcommand{\al}{\alpha}
\newcommand{\be}{\beta}
\newcommand{\ga}{\gamma}

\newcommand{\de}{\delta}

\newcommand{\un}[1]{{\underline{#1}} }

\newcommand{\tr}{\mathop{\rm tr}}

\newcommand{\mdeg}{\mathop{\rm mdeg}}
\newcommand{\Char}{\mathop{\rm char}}

\newcommand{\FF}{{\mathbb{F}}}   % base field
\newcommand{\NN}{{\mathbb{N}}}
\newcommand{\I}{\mathbb{I}} % i^2=-1
\newcommand{\symmmatr}[6]{\left(
\begin{array}{ccc}
#1 & #2 & #3\\ 
#2 & #4 & #5\\ 
#3 & #5 & #6 \\
\end{array}
\right)}

\newcommand{\mylabel}[1]{}

\newcommand{\mycomment}[1]{}
\begin{document}
\title[Minimal generating and separating sets for orthogonal invariants]{Minimal generating and separating sets for $O(3)$-invariants of several matrices}

\thanks{This research was supported by FAPESP 2016/00541-0 and FAEPEX 2449/18}
%\thanks{This research was supported by RFBR 16-31-60111 (mol\_a\_dk).}

\author{Ronaldo Jos\'e Sousa Ferreira}
\address{Ronaldo Jos\'e Sousa Ferreira\\ 
Federal University of Maranh\~ao, Rua Santa Clara, 2010,
65940-000 Graja\'u, MA, Brazil}
\email{ronaldoj.sf@hotmail.com (Ronaldo Jos\'e Sousa Ferreira)}

\author{Artem Lopatin}
\address{Artem Lopatin\\ 
Universidade Estadual de Campinas (UNICAMP), 651 Sergio Buarque de Holanda, 13083-859 Campinas, SP, Brazil}
\email{dr.artem.lopatin@gmail.com (Artem Lopatin)}
%\address{Artem Lopatin\\
%Sobolev Institute of Mathematics, Omsk Branch, SB RAS, Omsk, Russia}
%\email{artem\_lopatin@yahoo.com}

\begin{abstract} 
Given an algebra  $\FF[H]^G$ of polynomial invariants of an action of the group $G$ over the vector space $H$, 
a subset $S$ of $\FF[H]^G$ is called separating if $S$ separates all orbits that can be separated by $\FF[H]^G$. 
A minimal separating set is found for some algebras of matrix invariants of several matrices over an infinite field of arbitrary characteristic different from two in case of the orthogonal group. Namely, we consider the following cases:
\begin{enumerate}
\item[$\bullet$] $GL(3)$-invariants of two matrices;

\item[$\bullet$]  $O(3)$-invariants of $d>0$ skew-symmetric matrices; 

\item[$\bullet$]  $O(4)$-invariants of two skew-symmetric matrices;

\item[$\bullet$]  $O(3)$-invariants of two symmetric matrices.
\end{enumerate}
A minimal generating set is also  given for the algebra of orthogonal invariants of three $3\times 3$ symmetric matrices.

\noindent{\bf Keywords: } invariant theory, matrix invariants, classical linear groups,  separating invariants, generators, positive characteristic.

\noindent{\bf 2010 MSC: } 16R30; 15B10; 13A50.
\end{abstract}

\maketitle

%========================================================
%S1======================================================
\section{Introduction}\label{section_intro}

\subsection{Definitions} All vector spaces, algebras, and modules are over an infinite field $\FF$ of an arbitrary characteristic $p=\Char{\FF}\geq0$, unless otherwise stated.  By an algebra we always mean an associative algebra.

Given $n>1$ and $d\geq 1$,  we consider  the polynomial algebras 
$$\begin{array}{lcll}
R & = &\FF[x_{ij}(k) \;| \;1\leq i,j\leq n,& 1\leq k\leq d]; \\
R_{+} & = &\FF[x_{ij}(k)\;|\;1\leq j\leq i\leq n,& 1\leq k\leq d]; \\
R_{-} & = &\FF[x_{ij}(k)\;|\;1\leq j<i\leq n,& 1\leq k\leq d]. \\
\end{array}
$$
together with $n\times n$ {\it generic} matrices $X_k$,  {\it symmetric  generic} matrices $Y_k$ and {\it skew-symmetric generic} matrices  $Z_k$:
$$(X_k)_{ij}=x_{ij}(k),\qquad
(Y_k)_{ij}=
\left\{ 
\begin{array}{rc}
x_{ij}(k),& \text{ if } i\geq j\\
x_{ji}(k),& \text{ if } i< j\\
\end{array}
\right.,\qquad
(Z_k)_{ij}=
\left\{ 
\begin{array}{rc}
x_{ij}(k),& \text{ if } i> j\\
0 ,& \text{ if } i= j\\
-x_{ji}(k),& \text{ if } i< j\\
\end{array}
\right..
$$
Here $(A)_{ij}$ stands for the $(i,j)^{\rm th}$ entry of a matrix $A$.
The $t^{\rm th}$ coefficient of the characteristic polynomial of an $n\times n$ matrix $A$ is denoted by $\si_t(A)$. As an example, $\tr(A)=\si_1(A)$ and $\det(A)=\si_n(A)$.  Denote by $M(n)$ the space of all $n\times n$ matrices over $\FF$, $S_{+}(n)=\{A\in M(n)\,|\,A^T=A\}$, $S_{-}(n)=\{A\in M(n)\,|\,A^T=-A\}$ and $O(n)=\{A\in M(n)\,|\,AA^T=I_n\}$. Consider the algebras of {\it matrix invariants} $R^{GL(n)}$, $R^{O(n)}$, $R_{+}^{O(n)}$, $R_{-}^{O(n)}$, respectively, that are generated by  $\si_t(b)$, where $1\leq t\leq n$ and $b$ ranges over all monomials in \begin{enumerate}
\item[$\bullet$]  $X_1,\ldots,X_d$ (see~\cite{Sibirskii_1968}, \cite{Procesi76}, \cite{Donkin92a}),

\item[$\bullet$]  $X_1,\ldots,X_d,X_1^T,\ldots,X_d^{T}$ (see~\cite{Procesi76},~\cite{Zubkov99}), where $p\neq2$, 

\item[$\bullet$] $Y_1,\ldots,Y_d$ (see~\cite{ZubkovI} or~\cite{Lopatin_so_inv}), where $p\neq2$, 

\item[$\bullet$] $Z_1,\ldots,Z_d$ (see~\cite{ZubkovI} or~\cite{Lopatin_so_inv}), where $p\neq2$, 
\end{enumerate}
respectively.  Note that in case $p=0$ or $p>n$ the algebras of invariants considered above are generated by $\tr(b)$, where $b$ is the same as above. 
In what follows, whenever we consider the orthogonal group $O(n)$ or algebras $R^{O(n)}$, $R_{+}^{O(n)}$, $R_{-}^{O(n)}$, we assume that $p\neq2$. The ideal of relations between the generators of $R^{GL(n)}$ was described in~\cite{Razmyslov74, Procesi76, Zubkov96}. In case $p=0$ relations  between generators of  $R^{O(n)}$  were computed in~\cite{Procesi76} and in case $p\neq2$ relations between generators of matrix $O(n)$-invariants were obtained in~\cite{Lopatin_Orel} and~\cite{Lopatin_Ofree}.

The elements of $R$, $R_{+}$, $R_{-}$, respectively, can be interpreted as polynomial functions from 
 \begin{enumerate}
 \item[$\bullet$] $H=M(n)\oplus \cdots\oplus M(n)$, 
 
 \item[$\bullet$] $H_{+}=S_{+}(n)\oplus \cdots\oplus S_{+}(n)$,
  
 \item[$\bullet$] $H_{-}=S_{-}(n)\oplus \cdots\oplus S_{-}(n)$,
 \end{enumerate}
 respectively, to $\FF$ as follows: $x_{ij}(k)$ sends $u=(A_1,\ldots,A_d)\in H$ to $(A_k)_{i,j}$. We can consider $H$ as $GL(n)$-module by the formula: $g\cdot v  = (g A_1 g^{-1}, \ldots, g A_d g^{-1})$ for $g\in GL(n)$ and $v=(A_1,\ldots,A_d)\in H$. Then $H_{+}$ and $H_{-}$ are $O(n)$-modules. 
 
 Assume that $(G,A,V)$ is one of the following triples: $(GL(n),R,H)$, $(O(n),R,H)$, $(O(n),R_{+},H_{+})$, $(O(n),R_{-},H_{-})$. Then
 $$A^G=\{f\in A\,|\,f(g\cdot v)=f(v) \text{ for all }g\in G,\; v\in V\}$$
 (see the papers above mentioned, where the generators for the algebras of invariants were found).
 
 The notion of separating invariants was  introduced in 2002 by Derksen and Kemper~\cite{DerksenKemper_book} as a weaker concept than generating invariants.   Given a subset $S$ of $A^G$, we say that elements $u,v$ of $V$ {\it are separated by $S$} if  exists an invariant $f\in S$ with $f(u)\neq f(v)$. If  $u,v\in V$ are separated by $A^G$, then we simply say that they {\it are separated}. A subset $S\subset A^G$ of the invariant ring is called {\it separating} if for any $u, v$ from $V$ that are separated we have that they are separated
 by $S$. Separating sets over finite fields were studied in a recent paper~\cite{Kemper_Lopatin_Reimers_2}. 
 
 It follows from more general result of Domokos~\cite{Domokos_2007} and Draisma, Kemper, Wehlau~\cite{DraismaKemperWehlau_2008} that for any $n>1$ there exists $C(n)$, which does not depend on $d$, such that the set of all elements of $A^G$ of degree less than $C(n)$ is separating for all $d$. On the other hand, in case $0<p\leq n$ a similar statement is not valid for generating systems for $R^{GL(n)}$  (see~\cite{DKZ_2002}) and $R^{O(n)}$  (see~\cite{Lopatin_IndOrt}).
 
In~\cite{Lopatin_separating2x2} it was established that the set  
$$\begin{array}{cl}
\tr(X_i), \det(X_i),& 1\leq i\leq d,\\
\tr(X_i X_j), & 1\leq i<j\leq d,\\
\tr(X_i X_j X_k),&  1\leq i<j<k\leq d.\\
\end{array}
$$
is a minimal (by inclusion) separating set for the algebra of matrix invariants $R^{GL(2)}$ for any $d\geq1$.  The case of three nilpotent $3\times 3$ matrices over an algebraically closed field of zero characteristic was considered in~\cite{Cavalcante_Lopatin_1}.  A minimal separating set for the algebra $R^{SL(2)\times SL(2)}$ of semi-invariants of $2\times 2$ matrices over an arbitrary algebraically closed field was explicitly described in~\cite{Domokos20, Domokos20Add}. 

In this paper we will establish minimal (by inclusion) separating sets for the following algebras of invariants:
\begin{enumerate}
 \item[$\bullet$]  $R^{GL(3)}$ for $d=2$ (see~Theorem~\ref{theo1xx}), namely, 
 $$\begin{array}{c}
\si_t(X_i),\; i=1,2,\;t=1,2,3;\\
\tr(X_1 X_2),\;  \tr(X_1^2 X_2),\;\tr(X_1 X_2^2),\; \tr(X_1^2 X_2^2),\;
\tr(X_1^2 X_2^2 X_1 X_2);\\
\end{array} 
$$

\item[$\bullet$] $R^{O(3)}_{-}$ for all $d>0$, where $p\neq 2$ (see Theorem~\ref{theo3xx}), namely,
$$\begin{array}{c}
\si_2(Z_i);\;\tr(Z_i Z_j),\; i<j;\; \tr(Z_i Z_j Z_k),\;  i<j<k,\\
\end{array}
$$
where $1\leq i,j,k\leq d$;

\item[$\bullet$] $R^{O(4)}_{-}$ for $d=2$, where $p\neq 2$  (see Theorem~\ref{theo4xx}), namely,
$$\begin{array}{c}
\si_2(Z_i),\; \det(Z_i),\;i=1,2;\\
\tr(Z_1 Z_2),\;\;  \si_2(Z_1 Z_2),\; \tr(Z_1^2 Z_2^2),\; \tr(Z_1^3 Z_2),\; \tr(Z_1 Z_2^3);\\
\end{array}
$$

\item[$\bullet$] $R_{+}^{O(3)}$ for $d=2$, where $p\neq 2$  (see Theorem~\ref{theo-O3-symm2}), namely,
$$\begin{array}{c}
\si_t(Y_i),\; i=1,2,\;t=1,2,3;\;\; \tr(Y_1 Y_2),\;  \tr(Y_1^2 Y_2),\;\tr(Y_1 Y_2^2),\; \tr(Y_1^2 Y_2^2).\\
\end{array}
$$
We will establish that the last set is a minimal generating set for $R_{+}^{O(3)}$ for $d=2$.
\end{enumerate}

We will also construct a minimal generating set for  $R_{+}^{O(3)}$ for $d=3$  (see Theorem~\ref{theo-O3-symm3}), namely,
$$\begin{array}{c}
\si_t(Y_i),\; i,t\in\{1,2,3\};\;\;
\tr(Y_i Y_j),\;  \tr(Y_i^2 Y_j),\;\tr(Y_i Y_j^2),\; \tr(Y_i^2 Y_j^2),\; 1\leq i<j\leq 3;\\
\tr(Y_1 Y_2 Y_3),\\
\tr(Y_1^2 Y_2 Y_3),\;\tr(Y_2^2 Y_1 Y_3),\;\tr(Y_3^2 Y_1 Y_2),\;\;
\tr(Y_1^2 Y_2^2 Y_3),\;\tr(Y_1^2 Y_3^2 Y_2),\;\tr(Y_2^2 Y_3^2 Y_1).\\
\end{array}
$$

Note that over a field of real numbers a minimal generating set for $R_{+}^{O(3)}$ for each $d>0$ was constructed by Spencer and Rivlin in series of papers~\cite{Spencer1,Spencer2,Spencer3,Spencer4} (see also~\cite{Spencer_book}).

%Let us remark that Jing~\cite{Jing_2015} established that in case $\FF=\CC$ the algebra of invariants $R^{O(n)}$ separates the orbits of the action of $O(n)$ on $H$. Hence the same result is obviously holds for the pairs $(R_{+}^{O(n)}, H_{+})$ and $(R_{-}^{O(n)}, H_{-})$.
% !!!??? Chto s Jingom delat' ??

%=========================================================
\subsection{Notations}
\label{section_notations}

%If for $A,B\in M(n)$ there exists $g\in GL(n)$ such that $gAg^{-1}=B$, then we write $A\sim B$. 
For a monomial $c\in R$ denote by $\deg{c}$ its {\it degree} and by $\mdeg{c}$ its {\it multidegree}, i.e., $\mdeg{c}=(t_1,\ldots,t_d)$, where $t_k$ is the total degree of the monomial $c$ in $x_{ij}(k)$, $1\leq i,j\leq n$, and $\deg{c}=t_1+\cdots+t_d$. Obviously, the algebras $R^{GL(n)}$, $R^{O(n)}$, $R_{+}^{O(n)}$, $R_{-}^{O(n)}$ have $\NN$-grading by degrees and $\NN^d$-grading by multidegrees. 

Denote by $E_{ij}$ the matrix such that the $(i,j)^{\rm th}$ entry is equal to one and the rest of entries are zeros. 
%The diagonal matrix with elements $a_1,\ldots,a_n$ we denote by $\diag(a_1,\ldots,a_n)$. 
For short, we write $J_3$ for $E_{12} + E_{23}$.

%=========================================================
%=========================================================
\section{Decomposable invariants}\label{section_dec}
In this section we assume $n=3$ and $d>0$. We consider some fact about decomposable invariants that we are going to apply later.  

Assume that a triple $(G,A,V)$ is the same as in Section~\ref{section_intro}. We say that an $\NN$-homogeneous invariant $f\in A^G$ is {\it decomposable} and write $f\equiv0$ if $f$ is a polynomial in $\NN$-homogeneous invariants of $A^G$ of strictly lower degree. If $f$ is not decomposable, then we say that $f$ is {\it indecomposable} and write $f\not\equiv0$. In case $f-h\equiv0$ we write $f\equiv h$. Denote by $P_{\FF}$ the ring of polynomials in $Y_1,\ldots,Y_d$ without free terms and by $P$ the set of (non-empty) products of $Y_1,\ldots,Y_d$ and set  $P_{1}=P\sqcup \{I_3\}$.

Consider the surjective homomorphism $\Psi: R^{GL(n)}\to R_{+}^{O(n)}$ defined by $x_{ij}(k)\to x_{ij}(k)$ in case $i\geq j$ and by $x_{ij}(k)\to x_{ji}(k)$ otherwise. Note that the image of $\tr(X_{i_1}\cdots X_{i_k})$ with respect to $\Psi$ is $\tr(Y_{i_1}\cdots Y_{i_k})$.

\begin{lemma}\label{lemma_dec}
Assume $p\neq 2$, $x,y\in P$ and $q\in P_{\FF}$. Then the next formulas hold in $R_{+}^{O(3)}$:
\begin{enumerate}
\item[(a)] $\tr(x y x^2 q)\equiv -\tr(x^2 y x q)$;

\item[(b)] $\tr(Y_1^2 Y_2^i Y_1 Y_3^j)\equiv 0$ for $i,j=1,2$;

\item[(c)] $\tr(Y_1^2 Y_2^2 Y_1 Y_2)\equiv 0$;

\item[(d)] $\tr(Y_1^2 Y_2^2 Y_3^2)\equiv 0$ if $p\neq 3$;

\item[(e)] $\tr(y^2 x^2 y x q)\equiv -\tr(x^2 y^2 x y q)$; 

\item[(f)] $\tr(Y_1^2 Y_2^2 Y_1 Y_2 Y_3^i)\equiv 0$ for $i=1,2$;

%\item[(g)] $\tr(Y_1^2 Y_2^2 Y_1 Y_3^2)\equiv 0$.

\end{enumerate}
\end{lemma}
\begin{proof}
\noindent {\bf (a)} Applying the homomorphism $\Psi$ to formula~(20) of~\cite{Lopatin_Comm1} we obtain the required. 

For the sake of completeness, we show that part (a) can also be proven by straightforward calculations. Namely, part (a) follows from the next equality, which holds for every $3\times 3$ matrices $A,B,C$ over any commutative ring:
$$\begin{array}{c}
\tr(A^2BAC) + \tr(ABA^2C) = \\
\tr(A) \biggl(- \tr(A^2CB) +  \tr(ABAC) + \tr(BA^2C)  - \tr(AB) \tr(AC) - \tr(B) \tr(A^2C)  \biggr) + \\
+\si_2(A) \biggl(  \tr(ACB) - \tr(BAC) + \tr(B) \tr(AC) \biggl)\\ 
- \det(A)\tr(BC)  + \tr(A^3C) \tr(B)  + \tr(A^2C) \tr(AB)  + \tr(AC) \tr(A^2B) .\\
\end{array}$$

\medskip
\noindent {\bf (b) } Applying the equality $\tr(A^T)=\tr(A)$ that holds for any matrix $A$  and part~(a) of the lemma we obtain 
$$ \tr(Y_1^2 Y_2^i Y_1 Y_3^j) = \tr(Y_3^j Y_1 Y_2^i Y_1^2) = \tr(Y_1 Y_2^i Y_1^2 Y_3^j) \equiv -\tr(Y_1^2 Y_2^i Y_1 Y_3^j)$$
in $R_{+}^{O(3)}$. Since $p\neq2$, the proof of part~(b) is completed. 

\medskip
\noindent {\bf (c) } Making the substitution $Y_3\to Y_2$ in part~(b) of this lemma, where $i=2$ and $j=1$, we obtain the required.

\medskip
\noindent {\bf (d) } Since $p\neq 3$, applying the homomorphism $\Psi$ to part~7 of Lemma~18
from~\cite{Lopatin_Comm2} we obtain
$$\tr(Y_1^2 Y_2^2 Y_3^2)\equiv -\tr(Y_1^2 Y_3^2 Y_2^2) = -\tr(Y_1^2 Y_2^2 Y_3^2).$$
Since we also have that $p\neq2$, the proof of part~(d) is completed.

\medskip
\noindent {\bf (e) } Formula~(14) of~\cite{Lopatin_Comm1}, which is valid for any $p$,  together with Lemma~3 of~\cite{Lopatin_Comm1} imply that the analogue of part~(e) of this lemma is valid for $R^{GL(3)}$. The application of homomorphism $\Psi$ concludes the proof of part~(e).

\medskip
\noindent {\bf (f) } Applying two times part~(a) of the lemma to 
$f=\tr(Y_3^i Y_2 Y_1 Y_2^2 Y_1^2)$ we obtain that 
$$f\equiv\tr(Y_3^i Y_2^2 Y_1^2 Y_2 Y_1)\equiv -\tr(Y_3^i Y_1^2 Y_2^2 Y_1 Y_2),$$
where the second equivalence follows from part~(e) of the lemma. On the other hand, $f=\tr(Y_3^i Y_1^2 Y_2^2 Y_1 Y_2)$ and the proof of part~(f) is completed. 

%\medskip
%\noindent {\bf (g) } Applying part~(a) of the lemma, we have
%$$\tr(Y_1^2 Y_2^2 Y_1 Y_3^2) = \tr(Y_3^2 Y_1 Y_2^2 Y_1^2)\equiv 
%-\tr(Y_3^2 Y_1^2 Y_2^2 Y_1) = - \tr(Y_1^2 Y_2^2 Y_1 Y_3^2).$$
%The proof of part~(g) is completed. 
\end{proof}

%=========================================================
%=========================================================
\section{Invariants of two $3\times 3$ matrices}
\label{section_newresult_2GL3}

\begin{theo}\label{theo1xx}
For $d=2$ the following set is a minimal separating set for the algebra $R^{GL(3)}$ of  $GL(3)$-invariants of two matrices: 
$$\begin{array}{c}
\tr(X_i),\;\si_2(X_i),\;\det(X_i),\; i=1,2,\\
\tr(X_1 X_2),\;  \tr(X_1^2 X_2),\;\tr(X_1 X_2^2),\; \tr(X_1^2 X_2^2),\\
\tr(X_1^2 X_2^2 X_1 X_2) .\\
\end{array}
$$
\end{theo}
\begin{proof}
Denote the set from the formulation of the theorem by $S$. It is well-known that $S$ generates $R^{GL(3)}$ (for example, see~\cite{Lopatin_Sib}). Thus $S$ is a separating set. To prove that $S$ is a minimal separating set we will show that for any  element $f\in S$ the set $S_0=S\backslash \{f\}$ is not separating. 

The case of $f$ from the list $\tr(X_i)$, $\si_2(X_i)$, $\det(X_i)$ $(i=1,2)$ is obvious. 

Assume $f=\tr(X_1 X_2)$. Then for $A_1=B_1=J_3$ (see Section~\ref{section_notations}), $A_2=E_{32}$, $B_2=E_{12}$ we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. 

Assume $f=\tr(X_1^2 X_2)$. Then for 
$$A_1=B_1=J_3,\qquad
A_2 = \left(
\begin{array}{ccc}
0 & 1 & 1\\ 
0 & 1 & -1\\ 
0 & 1 &-1 \\
\end{array}
\right),\quad 
B_2 = \left(
\begin{array}{ccc}
0 & 0 & 0\\ 
1 & 0 & 0\\ 
-1 & 0 &0 \\
\end{array}
\right)
$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. Similarly we consider the case of $f=\tr(X_1 X_2^2)$.

Assume $f=\tr(X_1^2 X_2^2)$. Then for
$$A_1=B_1=J_3,\qquad
A_2 = \left(
\begin{array}{ccc}
0 & 1 & 0\\ 
0 & -1 & 0\\ 
1 & 1 & 1 \\
\end{array}
\right),\quad 
B_2 = \left(
\begin{array}{ccc}
0 & 1 & 0\\ 
1 & 0 & 0\\ 
1 & 0 &0 \\
\end{array}
\right)
$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set.

Assume $f=\tr(X_1^2 X_2^2 X_1 X_2)$. Then for
$$A_1=B_1=J_3,\qquad
A_2 = \left(
\begin{array}{ccc}
1 & 0 & 0\\ 
1 & -1 & 0\\ 
1 & -1 & 0 \\
\end{array}
\right),\quad 
B_2 = \left(
\begin{array}{ccc}
-1 & 0 & 0\\ 
0 & 0 & 1\\ 
1 & 0 &1 \\
\end{array}
\right)
$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. The theorem is proven.
\end{proof}

%=======================================================
%=======================================================
\section{Orthogonal invariants of several $3\times 3$ skew-symmetric  matrices}
\label{section_newresult_skewO3}

\begin{lemma}\label{lemma31xx}
The set  
$$
\si_2(Z_1),\; \si_2(Z_2),\; \tr(Z_1 Z_2)
$$
is a minimal separating set for the algebra $I=R_{-}^{O(3)}$ for $d=2$ in case $p\neq 2$.
\end{lemma}
\begin{proof}
Denote the set from the formulation of the lemma by $S$. The set $S$ generates the algebra of invariants $I$ (for example, see Theorem 1.4 of~\cite{Lopatin_Oskew}). Thus $S$ is a separating set. To prove that $S$ is a minimal separating set we will show that for any  element $f\in S$ the set $S_0=S\backslash \{f\}$ is not separating. 

The case of $f=\si_2(Z_i)$ ($i=1,2$) is obvious.

Assume $f=\tr(Z_1 Z_2)$. Then for $A_1=A_2=B_1=-B_2=E_{12}-E_{21}$ we have that $(A_1,A_2)$ and $(B_1,B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1,B_2)$. Thus $S_0$ is not a separating set. The lemma is proven.
\end{proof}

\begin{theo}\label{theo3xx}
Assume $d>0$ and  $p\neq 2$. Then the set  
$$\begin{array}{c}
\si_2(Z_i),\;1\leq i\leq d,\\
\tr(Z_i Z_j),\;  1\leq i<j\leq d,\\
\tr(Z_i Z_j Z_k),\;  1\leq i<j<k\leq d,\\
\end{array}
$$
is a minimal separating set for the algebra $R_{-}^{O(3)}$ of $O(3)$-invariants of $d$ skew-symmetric matrices.
\end{theo}
\begin{proof} Denote the set from the formulation of the theorem by $S$. By Theorem 1.4 of~\cite{Lopatin_Oskew} the set $S$ generates the algebra of invariants $R^{O(3)}_{-}$. Hence $S$ is a separating set for $R_{-}^{O(3)}$. Applying Lemma~\ref{lemma31xx} we obtain that to prove that $S$ is a minimal separating set it is enough to show that in case $d=3$ the set $S_0=S\backslash \{\tr(Z_1Z_2Z_3)\}$ is not separating. 

For $A_1=B_1=E_{12}-E_{21}$, $A_3=B_3=E_{13}-E_{31}$, and 
$$A_2 = \left(
\begin{array}{ccc}
0 & 0 & 1\\ 
0 & 0 & 1\\ 
-1 & -1 & 0 \\
\end{array}
\right),\quad 
B_2 = \left(
\begin{array}{ccc}
0 & 0 & 1\\ 
0 & 0 & -1\\ 
-1 & 1 &0 \\
\end{array}
\right)
$$
we have that $u=(A_1,A_2,A_3)$ and  $v=(B_1,B_2,B_3)$ are not separated by $S_0$, but $\tr(Z_1 Z_2 Z_3)$ separates $u$ and $v$ in case $p\neq2$. Thus $S_0$ is not a separating set.
The theorem is proven.
\end{proof}

%=======================================================
%=======================================================
\section{Orthogonal invariants of two $4\times 4$ skew-symmetric  matrices}
\label{section_newresult_2skewO4}

\begin{theo}\label{theo4xx}
Assume $d=2$ and $p\neq 2$. Then the set  
$$\begin{array}{c}
\si_2(Z_i),\; \det(Z_i),\;i=1,2,\\
\tr(Z_1 Z_2),\;  \si_2(Z_1 Z_2),\; \tr(Z_1^2 Z_2^2),\; \tr(Z_1^3 Z_2),\;\tr(Z_1 Z_2^3),\\
\end{array}
$$
is a minimal separating set for the algebra $R_{-}^{O(4)}$ of $O(4)$-invariants of two skew-symmetric matrices.
\end{theo}
\begin{proof} Assume $d=2$. Denote the set from the formulation of the theorem by $S$. By Theorem~1.1 of~\cite{Lopatin_mgs_skewO4},  the set $S$ generates the algebra of invariants $R^{O(4)}_{-}$. Hence $S$ is a separating set for $R_{-}^{O(4)}$. To prove that $S$ is a minimal separating set we will show that for any  element $f\in S$ the set $S_0=S\backslash \{f\}$ is not separating.

If $f$ is $\si_2(Z_1)$ or $\tr(Z_1 Z_2)$, then $S_0$ is not separating by Lemma~\ref{lemma31xx}. 

Assume $f=\det(Z_1)$. For 
$$A_1 = \left(
\begin{array}{cccc}
 0 & 0 & 0 & 1\\ 
 0 & 0 & 1 & 0\\
 0 &-1 & 0 & 0\\ 
-1 & 0 & 0 & 0\\ 
\end{array}
\right),\qquad
B_1 = \left(
\begin{array}{cccc}
 0 & 1 & 0 & 1\\ 
-1 & 0 & 0 & 0\\
 0 & 0 & 0 & 0\\ 
-1 & 0 & 0 & 0\\ 
\end{array}
\right),
$$
$A_2 = B_2 = 0$ we have that $u=(A_1,A_2)$ and  $v=(B_1,B_2)$ are not separated by $S_0$, but $\det(Z_1)$ separates $u$ and $v$. Thus $S_0$ is not a separating set.

Assume $f=\si_2(Z_1 Z_2)$. For 
$$A_1 = \left(
\begin{array}{cccc}
 0 & 1 & 0 & 0\\ 
-1 & 0 & 0 & 0\\
 0 & 0 & 0 &-1\\ 
 0 & 0 & 1 & 0\\ 
\end{array}
\right),\quad
A_2 = B_2 = \left(
\begin{array}{cccc}
 0 & 0 & 1 & 0\\ 
 0 & 0 & 0 & 1\\
-1 & 0 & 0 & 0\\ 
 0 &-1 & 0 & 0\\ 
\end{array}
\right),
\quad
B_1 = \left(
\begin{array}{cccc}
 0 &-1 & 0 & 0\\ 
 1 & 0 & 0 & 0\\
 0 & 0 & 0 &-1\\ 
 0 & 0 & 1 & 0\\ 
\end{array}
\right)
$$
we have that $u=(A_1,A_2)$ and  $v=(B_1,B_2)$ are not separated by $S_0$, but $\si_2(Z_1 Z_2)$ separates $u$ and $v$. Thus $S_0$ is not a separating set.

Assume $f=\tr(Z_1^2 Z_2^2)$. For 
$$A_1 = \left(
\begin{array}{cccc}
 0 & 1 & 1 & 1\\ 
-1 & 0 &-1 &-2\\
-1 & 1 & 0 &-1\\ 
-1 & 2 & 1 & 0\\ 
\end{array}
\right),\qquad
A_2 = \left(
\begin{array}{cccc}
 0 &-1 & 0 & 1\\ 
 1 & 0 & 0 & 0\\
 0 & 0 & 0 & 0\\ 
-1 & 0 & 0 & 0\\ 
\end{array}
\right),$$
$$B_1 = \left(
\begin{array}{cccc}
 0 &-2 & 0 &-1\\ 
 2 & 0 & 0 & 2\\
 0 & 0 & 0 & 0\\ 
 1 &-2 & 0 & 0\\ 
\end{array}
\right),
\qquad
B_2 = \left(
\begin{array}{cccc}
 0 & 0 & 1 & 0\\ 
 0 & 0 & 0 & 0\\
-1 & 0 & 0 & 1\\ 
 0 & 0 &-1 & 0\\ 
\end{array}
\right)
$$
we have that $u=(A_1,A_2)$ and  $v=(B_1,B_2)$ are not separated by $S_0$, but $\tr(Z_1^2 Z_2^2)$ separates $u$ and $v$. Thus $S_0$ is not a separating set.

Assume $f=\tr(Z_1 Z_2^3)$. For 
$$A_1 = B_1 = \left(
\begin{array}{cccc}
 0 & 0 & 1 & 1\\ 
 0 & 0 & 1 & 1\\
-1 &-1 & 0 & 0\\ 
-1 &-1 & 0 & 0\\ 
\end{array}
\right),\;
A_2 = \left(
\begin{array}{cccc}
 0 & 1 & 0 & 1\\ 
-1 & 0 & 0 &-1\\
 0 & 0 & 0 & 1\\ 
-1 & 1 &-1 & 0\\ 
\end{array}
\right),\;
B_2 = \left(
\begin{array}{cccc}
 0 & 1 & 0 &-1\\ 
-1 & 0 & 0 & 1\\
 0 & 0 & 0 & 1\\ 
 1 &-1 &-1 & 0\\ 
\end{array}
\right)
$$
we have that $u=(A_1,A_2)$ and  $v=(B_1,B_2)$ are not separated by $S_0$, but $\tr(Z_1 Z_2^3)$ separates $u$ and $v$. Thus $S_0$ is not a separating set. The case of $f=\tr(Z_1^3 Z_2)$ is similar. The theorem is proven.
\end{proof}

%=========================================================
%=========================================================
\section{Orthogonal invariants of two $3\times 3$ symmetric matrices}
 
\begin{theo}\label{theo-O3-symm2} Assume that $p\neq 2$ and $d=2$. Then the set
$$\begin{array}{c}
\tr(Y_i),\;\si_2(Y_i),\;\det(Y_i),\; i=1,2,\\
\tr(Y_1 Y_2),\;  \tr(Y_1^2 Y_2),\;\tr(Y_1 Y_2^2),\; \tr(Y_1^2 Y_2^2)\\
\end{array}
$$
is a minimal generating set and a minimal separating set for the algebra of $O(3)$-invariants $R_{+}^{O(3)}$ of two symmetric matrices.
\end{theo} 
\begin{proof}
Denote by $S$ the set from the formulation of the theorem and by $S_{X}$ the result of substitutions $Y_i\to X_i$ ($i=1,2$) in $S$. It is well-known that $S_X\cup\{\tr(X_1^2 X_2^2 X_1 X_2)\}$ generates $R^{GL(3)}$ in case $d=2$ (for example, see~\cite{Lopatin_Sib}). Considering the  surjective homomorphism  $\Psi$ from Section~\ref{section_dec}  we obtain that $S\cup f$  generates $R_{+}^{O(3)}$, where $f=\tr(Y_1^2 Y_2^2 Y_1 Y_2)$. Since $f$ is decomposable in $R_{+}^{O(3)}$ by part~(b) of Lemma~\ref{lemma_dec}, then $S$ generates $R_{+}^{O(3)}$. Thus $S$ is a separating set. To prove that $S$ is a minimal separating set we will show that for any  element $f\in S$ the set $S_0=S\backslash \{f\}$ is not separating. 

Assume $f=\tr(Y_1 Y_2)$. Then for 
$$A_1=B_1=\symmmatr{0}{0}{1}{1}{0}{0}, \quad A_2=  \symmmatr{0}{0}{0}{0}{1}{0}, \quad B_2=\symmmatr{1}{0}{0}{-1}{0}{0}$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. 

Assume $f=\tr(Y_1^2 Y_2)$. Then for 
$$A_1=B_1=\symmmatr{0}{0}{0}{0}{1}{0}, \quad A_2=  \symmmatr{0}{1}{0}{0}{0}{0}, \quad B_2=\symmmatr{1}{0}{0}{-1}{0}{0}$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. 
The case of $f=\tr(Y_1 Y_2^2)$ is similar. 

Assume $f=\tr(Y_1^2 Y_2^2)$. Then for 
$$A_1=\symmmatr{0}{1}{0}{0}{0}{0}, \quad A_2=B_1=  \symmmatr{0}{0}{0}{0}{1}{0}, \quad B_2=\symmmatr{0}{0}{0}{1}{0}{-1}$$
we have that $(A_1,A_2)$ and  $(B_1, B_2)$ are not separated by $S_0$, but $f$ separates $(A_1,A_2)$ and $(B_1, B_2)$. Thus $S_0$ is not a separating set. 

The cases of $f=\si_k(Y_i)$ ($i=1,2$, $k=1,2,3$) are trivial. For the sake of completeness, we point out that in case $f=\det(Y_1)$ we consider
$$A_1=-B_1=\symmmatr{1}{0}{0}{1}{0}{-2},\quad A_2=B_2=0.$$

Therefore,  $S$ is a minimal separating set. This result together with the fact that  $S$ is a generating set imlply that $S$ is a minimal generating set for $R_{+}^{O(3)}$. 
\end{proof}

%=========================================================
%=========================================================
\section{Orthogonal invariants of three $3\times 3$ symmetric matrices}

\begin{theo}\label{theo-O3-symm3}
For $d=3$ consider the following set $S$:
$$\begin{array}{c}
\tr(Y_i),\;\si_2(Y_i),\;\det(Y_i),\; i=1,2,3,\\
\tr(Y_i Y_j),\;  \tr(Y_i^2 Y_j),\;\tr(Y_i Y_j^2),\; \tr(Y_i^2 Y_j^2),\; 1\leq i<j\leq 3,\\
\tr(Y_1 Y_2 Y_3),\\
\tr(Y_1^2 Y_2 Y_3),\;\tr(Y_2^2 Y_1 Y_3),\;\tr(Y_3^2 Y_1 Y_2),\\
\tr(Y_1^2 Y_2^2 Y_3),\;\tr(Y_1^2 Y_3^2 Y_2),\;\tr(Y_2^2 Y_3^2 Y_1).\\
\end{array}
$$
Then the set
\begin{enumerate}
\item[$\bullet$] $S$, if $p\neq 2,3$,

\item[$\bullet$] $S\cup \{\tr(Y_1^2 Y_2^2 Y_3^2)\}$, if $p=3$,
\end{enumerate}
is a minimal generating set for the algebra of $O(3)$-invariants $R_{+}^{O(3)}$ of three symmetric matrices.
\end{theo}
\begin{proof}Denote by $S_1$ the set from the formulation of the theorem. We split the proof into two parts, namely, at first we show that $S_1$ generates $R_{+}^{O(3)}$ and then we prove that $S_1$ is minimal.

\smallskip
\noindent{\bf (a)} We apply $\Psi$ to the (minimal) generating set for $R^{GL(3)}$ from Theorem~1 of~\cite{Lopatin_Sib} and obtain that $R_{+}^{O(3)}$ is generated by $S\cup G_1$ in case $p\neq 2,3$ and by $S\cup G_1\cup G_2$ in case $p=3$. Here $G_1$ is the set
$$\begin{array}{c}
f_{ij}=\tr(Y_i^2 Y_j^2 Y_i Y_j),\; i<j;\;\;\; \tr(Y_1 Y_3 Y_2); \;\;\; h_1=\tr(Y_1^2 Y_2^2 Y_3^2); \\
\tr(Y_i^2 Y_j Y_k),\; j>k;\;\;\; \tr(Y_i^2 Y_j^2 Y_k),\; i>j; \\
r_{ijk}=\tr(Y_i^2 Y_j Y_i Y_k),\; j<k;\;\;\; s_{ijk}=\tr(Y_i^2 Y_j^2 Y_i Y_k),\; 
\end{array}
$$
where $1\leq i,j,k\leq 3$ are pairwise different, and $G_2$ is the set
$$\begin{array}{c}
h_2=\tr(Y_1^2 Y_3^2 Y_2^2),\\
a_{ijk}=\tr(Y_i Y_j^2 Y_k^2 Y_j Y_k),\;\;\; b_{ijk}=\tr(Y_i^2 Y_j^2 Y_i Y_k^2),\;\;\; c_{ijk}=\tr(Y_i^2 Y_j^2 Y_k^2 Y_j Y_k),\\
\end{array}
$$
where $j<k$, $1\leq i,j,k\leq 3$ are pairwise different. Parts~(b), (c), (f) of Lemma~\ref{lemma_dec} imply that $f_{ij}$, $r_{ijk}$, $s_{ijk}$, $a_{ijk}$, $b_{ijk}$, $c_{ijk}$ are decomposable in $R_{+}^{O(3)}$. It follows from part~(d) of  Lemma~\ref{lemma_dec} that $h_1$ and $h_2$ are decomposable in $R_{+}^{O(3)}$ in case $p\neq3$. Finally, the equalities $\tr(Y_1 Y_3 Y_2) = \tr(Y_1 Y_2 Y_3)\in S$,  $\tr(Y_i^2 Y_j Y_k) = \tr(Y_i^2 Y_k Y_j)$, $\tr(Y_i^2 Y_j^2 Y_k) = \tr(Y_j^2 Y_i^2 Y_k)$,  and $h_1=h_2$ imply that $S_1$ generates the algebra $R_{+}^{O(3)}$.

\medskip
\noindent{\bf (b)} Since all elements of $S$ have pairwise different multidegrees, to show that $S$ is a minimal generating set it is enough to establish that all elements of $S_1$ are indecomposable in $R_{+}^{O(3)}$. Then by Theorem~\ref{theo-O3-symm2} we can only verify the following claims:
\begin{enumerate}
\item[(1)] $f_1=\tr(Y_1 Y_2 Y_3)$ is indecomposable,

\item[(2)] $f_2=\tr(Y_1^2 Y_2 Y_3)$  is indecomposable,

\item[(3)] $f_3=\tr(Y_1^2 Y_2^2 Y_3)$ is indecomposable,

\item[(4)] $f_4=\tr(Y_1^2 Y_2^2 Y_3^3)$ is indecomposable in case $p=3$.
\end{enumerate}

Assume that $f_1$ is decomposable, i.e., $f_1$ is a polynomial in invariants of lower degree. Then it is easy to see that
\begin{eq}\label{eq1}
\tr(A_1 A_2 A_3)=0
\end{eq}%
for all nilpotent  matrices $A_1, A_2, A_3\in S_{+}(3)$ with entries in  $\FF$. Moreover, equality~(\ref{eq1}) is valid for any extension of $\FF$. In particular, we can assume that $\FF$ is algebraically closed. 

We will use the following symmetric nilpotent matrices as tests:
$$R_1 = \symmmatr{0}{1}{0}{0}{\I}{0},\; R_2 = \symmmatr{1}{1}{0}{-3}{2\I \sqrt{2}}{2},\; R_3 = \symmmatr{0}{-1}{0}{0}{\I}{0}$$
$$T_1 = \symmmatr{1}{\I}{0}{-1}{0}{0},\; 
T_2 =  \symmmatr{1}{-\I}{0}{-1}{0}{0},\;
T_3 =  \symmmatr{1}{0}{\I}{0}{0}{-1},\; 
$$
where $\I^2=-1$. Note that $T_i^2=0$ for $i=1,2,3$. 

We have $\tr(T_1 T_2 T_3)=2\neq0$; a contradiction to equality~(\ref{eq1}).

Assume that $f_2$ is decomposable. Then it is easy to verify that there exists $\al\in\FF$ such that
\begin{eq}\label{eq2}
\tr(A_1^2 A_2 A_3)= \al \tr(A_1 A_2) \tr(A_1 A_3)
\end{eq}%
for all nilpotent  matrices $A_1, A_2, A_3\in S_{+}(3)$ with entries in  $\FF$, which we assume to be algebraically closed. For $\un{A}=(A_1,A_2,A_3)=(T_1,T_2,T_3)$ equality~(\ref{eq2}) implies $\al=0$. Hence for $\un{A}=(R_1,R_2,T_1)$ equality~(\ref{eq2}) imply $1+(1-2\sqrt{2})\I=0$. It is easy to see that the obtained equality does not hold in case $p\neq2$.

Assume that $f_3$ is decomposable. Then it is easy to verify that there exist $\al,\be,\ga\in\FF$ such that
\begin{eq}\label{eq3}
\tr(A_1^2 A_2^2 A_3)= \al \tr(A_1 A_2) \tr(A_1 A_2 A_3) + 
\be \tr(A_1 A_3) \tr(A_1 A_2^2) +
\ga \tr(A_2 A_3) \tr(A_1^2 A_2)
\end{eq}%
for all nilpotent matrices $A_1, A_2, A_3\in S_{+}(3)$ with entries in  $\FF$, which we assume to be algebraically closed. Considering  $\un{A}=(T_1, T_2,T_3)$ in equality~(\ref{eq3}) we obtain $\al=0$. For $\un{A}=(T_1, R_1, T_2)$  equality~(\ref{eq3}) implies $\be=0$. If $\un{A}=(R_1, T_1, T_2)$ in equality~(\ref{eq3}), then we have $\ga=0$. Finally, for $\un{A}=(R_1, R_2, T_1)$ equality~(\ref{eq3}) imply $2(\I-1)(\sqrt{2}-1)=0$; a contradiction. 

Assume that $f_4$ is decomposable. Then it is easy to verify that there exist $\al_i,\be_i,\ga,\de\in\FF$ ($i=1,2,3$) such that  $\tr(A_1^2 A_2^2 A_3^3)=$ 
$$\begin{array}{c}
\al_1  \tr(A_1^2 A_2 A_3) \tr(A_2 A_3) + \al_2  \tr(A_2^2 A_1 A_3) \tr(A_1 A_3) +
   \al_3  \tr(A_3^2 A_1 A_2) \tr(A_1 A_2) + \\
  \be_1 \tr(A_1^2 A_3) \tr(A_2^2 A_3) + \be_2 \tr(A_1^2 A_2) \tr(A_3^2 A_2) +
  \be_3 \tr(A_2^2 A_1) \tr(A_3^2 A_1) +\\   
  \ga \tr(A_1 A_2 A_3)^2  + \de \tr(A_1 A_2) \tr(A_1 A_3) \tr(A_2 A_3)\\
\end{array}$$%
for all nilpotent matrices $A_1, A_2, A_3\in S_{+}(3)$ with entries in  $\FF$, which we assume to be algebraically closed. Making the substitution $\un{A}=(T_1,T_2,T_3)$ in the above equality we obtain $\de=-\ga$. Then,  consequently considering substitutions $\un{A}=(R_1, T_1, T_2)$, $\un{A}=(T_1,R_1, T_2)$ and $\un{A}=(T_1, T_2, R_1)$ we obtain that $\al_1=\al_2=\al_3=2\ga$. Similarly, substitutions $\un{A}=(R_1,R_2,T_1)$, $\un{A}=(R_1,T_1,R_2)$ and $\un{A}=(T_1,R_1,R_2)$ imply that $\be_1=\be_2=\be_3=-\ga$. Finally, applying the substitution $\un{A}=(R_1,R_2,R_3)$ we get $6 \ga = -1$, which is a contradiction in case $p=3$. Therefore, $S_1$ is a minimal generating set for $R^{O(3)}_{+}$.
\end{proof}

\end{document}